At 12-noon flight 237 leaves Chicago headed due south and traveling at a rate of r miles per hour. At 12-noon a second flight, flight 875, left Chicago from a different runway headed due east and traveling 40 miles per hour slower than flight 237. At 3:00pm the planes are 600 miles apart. How fast is flight 237 traveling?
Please assist with the set up for the equation.

Answer by rothauserc(4718)

you have a right triangle with the hypotenuse = 600 miles
use equation rate * time = distance
let r be the rate of flight 237 and r-40 is rate for flight 875
therefore, 3r is distance flown by flight 237 and 3(r-40) is the distance flight 875 flies
now we can set up the equation using Pythagorean Theorem
(3r)^2 + (3r-120)^2 = 600^2
9r^2 + 9r^2 -720r +14400 = 360000
18r^2 -720r -345600 = 0
r^2 -40r -19200 = 0
you can solve this using quadratic formula
r = (40 + square root(40^2 -4*(-19200))) / 2
r = (40 - square root(40^2 -4*(-19200))) / 2